# Adaptive Quantum Control

### v1

@@ IV Numerical Simulation
In this paper, we consider a two-level system, with dynamics given by the LGKS (1), with a single damped term, that is, $n=1$. The control system architecture is shown in Figure 1.
$F(\rho,\rho_{\rm{d}})$$\begin{array}[]{c}\text{Adaptive}\\ \text{Controller}\end{array}$ $\mathcal{L}\mathcal{G}\mathcal{K}\mathcal{S}$$\rho_{\rm{d}}$$e$$u$$\rho$
Figure 1: Adaptive control architecture to track desired density matrices in the LGKS system.
To simulate the LKGS equation, we set
@@
The initial density matrix $\rho(0)$ is assumed to be
$\rho_{0}$ $=\begin{bmatrix}0.4&0.1+0.3\imath\\ 0.1-0.3\imath&0.6\end{bmatrix}.$ (27)
Note that $\rho(0)$ is randomly generated such that ${\rm tr}\,\rho(0)=1$ and it is Hermitian. As shown in [moses2015creation] and [xu2021multicriticality], we consider the problem of moving the quantum state $\rho(0)$ to a low and a high desired entropy states $\rho_{\rm d},$ where the Von Neumann entropy of a quantum state $\rho$ is given by
@@
First, we consider the problem of driving the quantum state $\rho(t)$ to
@@
Note that $S(\rho_{\rm d})=0.1013,$ and thus, this desired state is a low-entropy state. Furthermore, note that $\rho_{\rm d},$ given by (29) is an equilibrium point of (1) in the case where $u(t)\equiv 1.$
With the desired density $\rho_{\rm d}$ given by (29) as the reference for the controller, we first tune the RCAC hyperparameters. In particular, in RCAC, we set $R_{z}=1,$ $R_{u}=1,$ and $\lambda=0.01$; and optimize $P_{0}$ and the filter coefficient $\beta,$ where $G_{\rm f}(s)=\dfrac{1}{s+\beta},$ as described below. We optimize only $P_{0}$ and the filter coefficient since it has been observed in several numerical experiments that the choice of $P_{0}$ and the filter coefficient has the most dominant impact on the closed-loop performance. For each pair of $(P_{0},\beta)\in\{(10^{-5},10^{-4},\ldots,10^{10})\times(0,1,2,5,100,2000)\},$ we apply the RCAC algorithm to update the adaptive PID controller to drive the quantum state from $\rho_{0}$, given by (27) to $\rho_{\rm d}$, given by (29). Each simulation is run for 200 seconds. At the end of each simulation, the hyperparameter cost
@@
is computed. Figure 2 shows the contour plot of $J_{h}$ for various values of $(P_{0},\beta).$ The tuning pair with the minimum value of $J_{h}$ is selected for all the numerical simulations in the rest of the paper. In particular, in all the examples, we set $P_{0}=10^{-3}\times I_{3},\beta=2000.$
Figure 2: Contour plot of $J_{h}$ for various values of $(P_{0},\beta).$ The pair $(P0,\beta)$ that minimizes $J_{h}$ is used in all numerical simulations.
IV-A Low-entropy Density
With the selected RCAC hyperparameters, Figure 3 shows the absolute value of the density error $e(t)$ between the system state and the desired low-entropy state on a logarithmic scale, the adaptive gains $k_{\rm p},$ $k_{\rm i},$ and $k_{\rm d},$ updated by RCAC, and the control $u$ given by adaptive controller. Figure 4 shows the real and imaginary part of the quantum state $\rho(t)$ and the Von Neumann entropy of the quantum state. Figure 5 shows the trajectory of the quantum state of the quantum system (1) on the Bloch sphere. Note that the quantum state $\rho$ satisfies $x=2\,\rm{Real}(\rho_{1,2}),$ $y=2\,\rm{Imag}(\rho_{1,2}),$ and $z=\rho_{1,1}-\rho_{2,2}$ on the Bloch sphere [wie2020two].
Figure 3: Low-entropy density tracking. The first subplot shows the absolute value of the density error $e(t)$ between the system state and the desired state on a logarithmic scale, the next three subplots show the adaptive gains $k_{\rm p},$ $k_{\rm i},$ and $k_{\rm d},$ updated by RCAC, and the last subplot shows the control $u$ given by adaptive controller.
Figure 4: Low-entropy density tracking. The first two subplots show the real and imaginary parts of the quantum state $\rho(t).$ Note that the desired state component is shown in dashes and the system response is shown in solid with the corresponding color. The third subplot shows the entropy of the quantum state.
Figure 5: Low-entropy density tracking. Trajectory of the quantum state of the quantum system (1) on the Bloch sphere.
IV-B High-entropy Density
Next, we consider the problem of driving the quantum state $\rho(t)$ to
@@ IV-B High-entropy Density
Note that $S(\rho_{\rm d})=0.6693,$ and thus, this desired state is a high-entropy state. Furthermore, note that $\rho_{\rm d},$ given by (31) is an equilibrium point of (1) in the case where $u(t)\equiv 10.$
@@
Figure 6 shows the absolute value of the density error $e(t)$ between the system state and the desired high-entropy state on a logarithmic scale, the adaptive gains $k_{\rm p},$ $k_{\rm i},$ and $k_{\rm d},$ updated by RCAC, and the control $u$ given by adaptive controller. Figure 7 shows the real and imaginary part of the quantum state $\rho(t)$ and the Von Neumann entropy of the quantum state. Figure 8 shows the trajectory of the quantum state of the quantum system (1) on the Bloch sphere.
The two examples described above suggest that the adaptive PID controller updated by the RCAC algorithm can effectively track an arbitrary desired density matrix $\rho_{\rm d}$ without requiring any model of the underlying dynamics of the density matrix.
@@
Figure 6: High-entropy density tracking. The first subplot shows the absolute value of the density error $e(t)$ between the system state and the desired state on a logarithmic scale, the next three subplots show the adaptive gains $k_{\rm p},$ $k_{\rm i},$ and $k_{\rm d},$ updated by RCAC, and the last subplot shows the control $u$ given by adaptive controller.
@@
Figure 7: High-entropy density tracking. The first two subplot shows the real and imaginary part of the quantum state $\rho(t).$ Note that the desired state component is shown in dashes and the system response is shown in solid with the corresponding color. The third subplot shows the entropy of the quantum state.
@@
Figure 8: High-entropy density tracking. Trajectory of the quantum state of the quantum system (1) on the Bloch sphere.
@@ V Conclusions
This paper presented an online learning-based, model-free adaptive controller for tracking a desired density matrix in a two-level Lindblad-Gorini-Kossakowski-Sudarshan quantum system. In particular, the adaptive controller uses the retrospective cost adaptive control algorithm. Since the classical RCAC was developed for a discrete-time system, the paper presented a continuous-time extension of RCAC. An adaptive PID controller structure was chosen to generate a control signal that preserves the LKGS solution’s geometric structure. Uhlmann-Jozsa fidelity was used as an error metric to drive the adaptive PID controller. Numerical simulations show that the proposed controller effectively tracks desired low and high-entropy states without requiring a model of the underlying system dynamics.
Future work will extend the technique to the case where the input is a Hermitian matrix and the density coherence is used as an error metric. This will allow higher-dimensional parameterization of the adaptive PID controller, which may yield better tracking performance.
[…]
@@ A Some Useful Facts
Let $A\in{\mathbb{C}}^{n\times n}.$ $A$ is a Hermitian matrix if $A=\overline{A^{\rm T}},$ where $\overline{A^{\rm T}}$ is the conjugate transpose of $A.$
@@
Let $A,B\in{\mathbb{C}}^{n\times n}$ be Hermitian. Then, $-\imath[A,B]$ is Hermitian.
@@
Note that $(-i[A,B])^{\rm H}=(-\imath)^{\rm H}[A,B]^{\rm H}=\imath\left(AB-BA\right)^{\rm H}=\imath(A^{\rm H}B^{\rm H}-B^{\rm H}A^{\rm H})=\imath(BA-AB)=-\imath(AB-BA)=-\imath[A,B]$. Thus, $-\imath[A,B]$ is Hermitian. ∎
@@
$1\geq\left({\rm tr}\,\left(\sqrt{\sqrt{\rho}\,\sigma\,\sqrt{\rho}}\right)\right)^{2}\in\mathbb{R}\geq 0$
@@
Since $\rho$ and $\sigma$ are PSD Hermitian. Then, $\rho^{{\rm H}}=\rho$, $\sigma^{{\rm H}}=\sigma$, $x^{{\rm H}}\rho x\geq 0$ and $x^{{\rm H}}\sigma x\geq 0$ for all $x\in\mathbb{C}.$ Furthermore, the unique square root of a PSD Hermitian matrix is also Hermitian and PSD [horn2012matrix, p 440]. This implies that, $\left(\sqrt{(}\rho)\sigma\sqrt{\rho}\right)^{{\rm H}}=\sqrt{\rho}^{{\rm H}}\left(\sqrt{(}\rho)\sigma\right)^{{\rm H}}=\sqrt{\rho}^{{\rm H}}\sigma^{{\rm H}}\sqrt{\rho}^{{\rm H}}=\sqrt{\rho}\sigma\sqrt{\rho}.$ Define, $y\stackrel{{\scriptstyle\triangle}}{{=}}\sqrt{\rho}x\in\mathbb{C}^{n}.$ Then, $x^{{\rm H}}\sqrt{\rho}\sigma\sqrt{\rho}x=y^{{\rm H}}\sigma y\geq 0,$ which implies that $\sqrt{\rho}\sigma\sqrt{\rho}\geq 0.$ Therefore, $\left(\sqrt{(}\rho)\sigma\sqrt{\rho}\right)$ is Hermitian, and $\sqrt{\left(\sqrt{(}\rho)\sigma\sqrt{\rho}\right)}$ is Hermitian and unique. From the Cauchy-Schwarz inequality and the cyclic property for traces, $|{\rm tr}\,(\sqrt{\sqrt{\rho}\sigma\sqrt{\rho}})|^{2}=|{\rm tr}\,(\sqrt{\sqrt{\rho}\sqrt{\rho}\sigma})|^{2}=|{\rm tr}\,(\sqrt{\rho\sigma})|^{2}\leq{\rm tr}\,(\rho){\rm tr}\,(\sigma)\leq 1$. Thus, $1\geq\left({\rm tr}\,\left(\sqrt{\sqrt{\rho}\,\rho_{\rm{d}}\,\sqrt{\rho}}\right)\right)^{2}\in\mathbb{R}\geq 0$

### v2

@@ IV Numerical Simulation
In this paper, we consider a two-level LGKS system (1) with a single dissipative channel, that is, $m=1$. The control system architecture is shown in Fig. 1. Note that, for conceptual clarity, the observer is included in the block diagram; however, in the numerical studies, we assume access to an estimate of the density matrix $\rho$ for feedback computation. In practice, the quantum state is not directly measurable in real time and must be reconstructed from measurement records using quantum filtering or observer-based techniques; a representative observer design is described in [qamar2019observer]. In this preliminary study, the observer dynamics and measurement backaction are not explicitly modeled, and the focus is on evaluating the closed-loop behavior of the adaptive controller under idealized assumptions about state access.
To simulate the LGKS equation, we set
@@
The initial density matrix is chosen as
$\rho_{0}$ $=\begin{bmatrix}0.4&0.1+0.3\imath\\ 0.1-0.3\imath&0.6\end{bmatrix}$, (27)
which is a randomly generated Hermitian, positive semidefinite matrix with unit trace. Following [22, 34], we consider both low-entropy and high-entropy target states $\rho_{\rm d}$, where the von Neumann entropy of a quantum state $\rho$ is defined as
@@
Achieving or maintaining low-entropy states is often a central objective in quantum information processing, since low entropy corresponds to high purity and coherence, which are essential resources for quantum computation, communication, and sensing. Because environmental interactions and decoherence naturally increase entropy, feedback control is required to counteract this drift and preserve useful quantum states. Conversely, the ability to track high-entropy target states is also relevant in many physical scenarios, since open quantum systems are often driven toward mixed or thermal equilibrium states by their environment.
IV-A Hyperparameter Tuning
First, we specify the low-entropy target state
@@ IV-A Hyperparameter Tuning
Note that $S(\rho_{\rm d})=0.1013,$ and thus, this desired state is a low-entropy state. Furthermore, $\rho_{\rm d},$ given by (29), is an equilibrium point of (1) in the case where $u(t)\equiv 1.$
With the desired density $\rho_{\rm d}$ given by (29) as the reference for the controller, we tune the RCAC hyperparameters. In particular, we set $R_{z}=1,$ $R_{u}=1,$ and $\lambda=0.01,$ and optimize $P_{0}$ and the filter coefficient $\beta,$ where $G_{\rm f}(s)=\dfrac{1}{s+\beta}.$ We restrict the optimization to $P_{0}$ and $\beta,$ as numerical experiments consistently indicate that these parameters are the most dominant factors in the closed-loop performance.
For each pair of $(P_{0},\beta)\in\{(10^{-5},10^{-4},\ldots,10^{10})\times(0,1,2,5,100,2000)\},$ the RCAC algorithm is applied to update the adaptive PID controller to drive the quantum state from $\rho_{0}$, given by (27), to $\rho_{\rm d}$, given by (29). Each simulation is run for 200 seconds. At the end of each simulation, the hyperparameter cost
@@
is computed. Figure 2 shows the contour plot of $J_{h}$ for various values of $(P_{0},\beta).$ The tuning pair with the minimum value of $J_{h}$ is selected for all numerical simulations presented in the remainder of the paper. In particular, in all the examples, we set $P_{0}=10^{-3}\times I_{3}$ and $\beta=2000.$ We emphasize that the hyperparameter tuning procedure outlined above is coarse and heuristic in nature, and is intended to identify a representative near-optimal parameter set for the examples considered here.
Figure 2: Contour plot of $J_{h}$ for various values of $(P_{0},\beta).$ The pair $(P_{0},\beta)$ that minimizes $J_{h}$ is used in all numerical simulations.
IV-B Low-Entropy Density Tracking
With the selected RCAC hyperparameters, Fig. 3 shows the absolute value of the density error $e(t)$ between the detected system state and the desired low-entropy state on a logarithmic scale, the adaptive gains $k_{\rm p},$ $k_{\rm i},$ and $k_{\rm d},$ updated by RCAC, and the control $u$ generated by the adaptive controller. Fig. 4 shows the real and imaginary parts of the quantum state $\rho(t)$ and the von Neumann entropy of the quantum state. Fig. 5 shows the trajectory of the quantum state of the LGKS system (1) on the Bloch sphere. Note that the quantum state $\rho$ satisfies $x=2\,\rm{Real}(\rho_{1,2}),$ $y=2\,\rm{Imag}(\rho_{1,2}),$ and $z=\rho_{1,1}-\rho_{2,2}$ on the Bloch sphere [33].
Figure 3: Low-entropy density tracking. The first subplot shows the absolute value of the density error $e(t)$ between the system state and the desired state on a logarithmic scale, the next three subplots show the adaptive gains $k_{\rm p},$ $k_{\rm i},$ and $k_{\rm d},$ updated by RCAC, and the last subplot shows the control $u$ generated by the adaptive controller.
Figure 4: Low-entropy density tracking. The first two subplots show the real and imaginary parts of the quantum state $\rho(t).$ The desired state components are shown with dashed lines, while the system response is shown with solid lines of the corresponding color. The third subplot shows the entropy of the quantum state.
Figure 5: Low-entropy density tracking. Trajectory of the quantum state of the LGKS system (1) on the Bloch sphere. Note that the trajectory remains inside the Bloch ball, consistent with mixed-state evolution.
IV-C High-Entropy Density Tracking
Next, we consider the problem of driving the quantum state $\rho(t)$ to the high-entropy target
@@ IV-C High-Entropy Density Tracking
Note that $S(\rho_{\rm d})=0.6693,$ and thus, this desired state is a high-entropy state. Furthermore, $\rho_{\rm d},$ given by (31), is an equilibrium point of (1) in the case where $u(t)\equiv 10.$
@@
Figure 6 shows the absolute value of the density error $e(t)$ between the detected system state and the desired high-entropy state on a logarithmic scale, the adaptive gains $k_{\rm p},$ $k_{\rm i},$ and $k_{\rm d},$ updated by RCAC, and the control input $u$ generated by the adaptive controller. Figure 7 shows the real and imaginary parts of the quantum state $\rho(t)$ and the von Neumann entropy of the quantum state. Figure 8 shows the trajectory of the quantum state of the LGKS system (1) on the Bloch sphere.
@@
Figure 6: High-entropy density tracking. The first subplot shows the absolute value of the density error $e(t)$ between the system state and the desired state on a logarithmic scale, the next three subplots show the adaptive gains $k_{\rm p},$ $k_{\rm i},$ and $k_{\rm d},$ updated by RCAC, and the last subplot shows the control $u$ generated by the adaptive controller.
@@
Figure 7: High-entropy density tracking. The first two subplots show the real and imaginary parts of the quantum state $\rho(t).$ The desired state components are shown with dashed lines, while the system response is shown with solid lines of the corresponding color. The third subplot shows the entropy of the quantum state.
@@
Figure 8: High-entropy density tracking. Trajectory of the quantum state of the LGKS system (1) on the Bloch sphere. Note that the trajectory remains inside the Bloch ball, consistent with mixed-state evolution.
IV-D Shot Noise
Finally, we revisit the low-entropy target $\rho_{\rm d}$ given by (29) and examine the effect of measurement noise on the feedback loop. Shot noise in the feedback detection is modeled as additive white noise, following [niebauer1991nonstationary]. Specifically, at each discrete sampling instant, a perturbation $N\in\mathbb{C}^{2\times 2}$ is drawn with entries distributed as $\mathcal{N}(0,\sigma^{2})$, and the noisy state is constructed as
$\rho_{\text{noisy}}=\frac{\rho+\tfrac{1}{2}(N+N^{\dagger})}{{\rm tr}\,\!\left(\rho+\tfrac{1}{2}(N+N^{\dagger})\right)}.$ (32)
Hermiticity is enforced by symmetrization, and physical validity is preserved by trace normalization, ensuring that $\rho_{\text{noisy}}$ remains a valid density matrix under white noise of strength $\sigma=5\times 10^{-3}$.
Remark IV.1.
Noise model and feedback-only corruption. The noisy state $\rho_{\text{noisy}}$ is used only in the feedback path for computing the fidelity-based error signal. The underlying system evolution is still governed by the deterministic LGKS dynamics driven by the true state $\rho(t)$. Thus, the injected noise models the corruption of the measurement or state estimate rather than physical noise acting directly on the quantum dynamics.
Figure 9 shows the absolute value of the density error $e(t)$ between the detected system state with shot noise and the desired state on a logarithmic scale, the adaptive gains $k_{\rm p},$ $k_{\rm i},$ and $k_{\rm d},$ updated by RCAC, and the control $u$ generated by the adaptive controller. Figure 10 shows the real and imaginary parts of the quantum state $\rho(t)$ and the von Neumann entropy of the quantum state. Figure 11 shows the trajectory of the quantum state of the LGKS system (1) on the Bloch sphere.
The three examples described above suggest that the adaptive PID controller updated by the RCAC algorithm can effectively track a desired density matrix $\rho_{\rm d}$ in the presence of moderate measurement noise, without requiring a model of the underlying LGKS dynamics in the control law.
Figure 9: Shot noise case tracking. The first subplot shows the absolute value of the density error $e(t)$ between the detected system state with shot noise and the desired state on a logarithmic scale, the next three subplots show the adaptive gains $k_{\rm p},$ $k_{\rm i},$ and $k_{\rm d},$ updated by RCAC, and the last subplot shows the control $u$ generated by the adaptive controller.
Figure 10: Shot noise case tracking. The first two subplots show the real and imaginary parts of the quantum state $\rho(t).$ The desired state components are shown with dashed lines, while the system response is shown with solid lines of the corresponding color. The third subplot shows the entropy of the quantum state.
Figure 11: Shot noise case tracking. Trajectory of the quantum state of the LGKS system (1) on the Bloch sphere. Note that the trajectory remains inside the Bloch ball, consistent with mixed-state evolution.
@@ V Conclusions
This paper presented an online learning-based adaptive control framework for tracking a desired density matrix in a two-level Lindblad–Gorini–Kossakowski–Sudarshan quantum system, in which the feedback control law does not require prior knowledge of the system Hamiltonian or dissipative operators. In particular, the adaptive controller is based on the retrospective cost adaptive control (RCAC) algorithm, and a continuous-time formulation is employed to align with the LGKS dynamics. An adaptive PID controller structure is used to generate the control signal while preserving the geometric structure of the LGKS evolution. Uhlmann–Jozsa fidelity is adopted as a scalar performance metric to drive the adaptation. Numerical simulations demonstrate effective tracking of both low-entropy and high-entropy target states under idealized state-access assumptions and in the presence of moderate measurement noise.
Future work will extend this framework to matrix-valued control inputs and alternative quantum error metrics, such as coherence-based measures, enabling richer controller parameterizations and evaluation on higher-dimensional quantum systems. Additional directions include incorporating stochastic measurement backaction, discrete-time feedback implementations, and theoretical analysis of closed-loop performance guarantees.
@@ A Some Useful Facts
Let $\in{\mathbb{C}}^{n\times n}. AA$ is a Hermitian matrix if $A=\overline{A^{\rm T}},$ where $\overline{A^{\rm T}}$ is the conjugate transpose of $A.$
@@
Let $A,B\in{\mathbb{C}}^{n\times n}$ be Hermitian. Then, $\imath[A,B]$ is Hermitian.
@@
Note that $(-i[A,B])^{\rm H}=(-\imath)^{\rm H}[A,B]^{\rm H}=\imath\left(AB-BA\right)^{\rm H}=\imath(A^{\rm H}B^{\rm H}-B^{\rm H}A^{\rm H})=\imath(BA-AB)=-\imath(AB-BA)=-\imath[A,B]$. Thus, $\imath[A,B]$ is Hermitian. ∎
@@
$1\geq\left({\rm tr}\,\left(\sqrt{\sqrt{\rho}\,\sigma\,\sqrt{\rho}}\right)\right)^{2}\in\mathbb{R}\geq 0.$
@@
Since $\rho$ and $\sigma$ are PSD Hermitian. Then, $\rho^{{\rm H}}=\rho$, $\sigma^{{\rm H}}=\sigma$, $x^{{\rm H}}\rho x\geq 0$ and $x^{{\rm H}}\sigma x\geq 0$ for all $x\in\mathbb{C}.$ Furthermore, the unique square root of a PSD Hermitian matrix is also Hermitian and PSD [13, p 440]. This implies that, $\left(\sqrt{(}\rho)\sigma\sqrt{\rho}\right)^{{\rm H}}=\sqrt{\rho}^{{\rm H}}\left(\sqrt{(}\rho)\sigma\right)^{{\rm H}}=\sqrt{\rho}^{{\rm H}}\sigma^{{\rm H}}\sqrt{\rho}^{{\rm H}}=\sqrt{\rho}\sigma\sqrt{\rho}.$ Define, $y\stackrel{{\scriptstyle\triangle}}{{=}}\sqrt{\rho}x\in\mathbb{C}^{n}.$ Then, $x^{{\rm H}}\sqrt{\rho}\sigma\sqrt{\rho}x=y^{{\rm H}}\sigma y\geq 0,$ which implies that $\sqrt{\rho}\sigma\sqrt{\rho}\geq 0.$ Therefore, $\left(\sqrt{(}\rho)\sigma\sqrt{\rho}\right)$ is Hermitian, and $\sqrt{\left(\sqrt{(}\rho)\sigma\sqrt{\rho}\right)}$ is Hermitian and unique. From the Cauchy-Schwarz inequality and the cyclic property for traces, $|{\rm tr}\,(\sqrt{\sqrt{\rho}\sigma\sqrt{\rho}})|^{2}=|{\rm tr}\,(\sqrt{\sqrt{\rho}\sqrt{\rho}\sigma})|^{2}=|{\rm tr}\,(\sqrt{\rho\sigma})|^{2}\leq{\rm tr}\,(\rho){\rm tr}\,(\sigma)\leq 1$. Thus, $1\geq\left({\rm tr}\,\left(\sqrt{\sqrt{\rho}\,\rho_{\rm{d}}\,\sqrt{\rho}}\right)\right)^{2}\in\mathbb{R}\geq 0.$